OpenStudy (anonymous):

Show me the process of elimination on this problem? 6x – 4y = 54 –9x + 2y = –69

OpenStudy (ajprincess):

Multiply the second eqn by 2. Vat do u get?

OpenStudy (ajprincess):

Vat do u get ven u multiply the second eqn by 2?

OpenStudy (aroub):

Okay, inorder to do the elimination process here as ajprincess mentioned above you multiply the second equation to cancel 4y: 6x – 4y = 54 -18x+4y=-138 -------------- -12x=-84 x=7 If you want to find y you substitute x by either of the equations: 6(7)-4y=54 -4y+42=54 -4y=12 y=-3 Check: 6(7)-4(-3)=54 42+12=54 54=54

OpenStudy (anonymous):

is this right??

OpenStudy (anonymous):

is this like the only thing you do

OpenStudy (aroub):

Yup :)